# Style-Adaptive Detection Transformer for Single-Source Domain Generalized Object Detection

### v1

@@ III-B Domain Style Adapter
Given the feature map of a specific layer $F_{d}\in\mathbb{R}^{(B\times C\times H\times W)}$ (where $d$ denotes the domain of the input image, either the source domain or the augmented domain), we first compute its statistical data, i.e., the mean $\mu_{d}$ and variance $\sigma_{d}$ along the channel dimension, as follows:
$\mu_{d}\left(F_{d}\right)=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}F_{d},$ (1)
$\sigma_{d}\left(F_{d}\right)=\sqrt{\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}\left(F_{d}-\mu_{d}\left(F_{d}\right)\right)^{2}+\epsilon},$ (2)
where $\mu_{d}$ and $\sigma_{d}\in\mathbb{R}^{(B\times C)}$, $\epsilon$ is set to $1\times 10^{-6}$ to avoid numerical computation resulting in zero. Since the network’s input consists of image pairs, we calculate the mean and variance for both the source and the augmented domain images, respectively.
Then, we store the means and variances for each domain using a memory module during each training iteration and utilize the Exponential Moving Average (EMA)[42] method to generate domain-specific style bases online, as detailed below:
$p_{d}^{\mu}=\alpha p_{d}^{\mu}+(1-\alpha)\mu_{d},$ (3)
$p_{d}^{\sigma}=\alpha p_{d}^{\sigma}+(1-\alpha)\sigma_{d},$ (4)
where $\alpha$ represents the momentum factor, which is normally close to 1. Both $p_{d}^{\mu}$ and $p_{d}^{\sigma}$ are initialized to zero. As the stored statistical data accumulate, the memory module can develop dataset-level style represents.
After the calculations described above, we obtain two style bases corresponding to the source and augmented domains. Subsequently, we use these bases to rectify the style of the input image, projecting the style representation of the target domain into the training domain. Specifically, we use the Wasserstein distance[18] to measure the distribution differences in the statistical data between the current image and the two stored bases to obtain the correlation coefficient, as follows:
$\lambda_{d}=\left\|\mu_{d}-p_{d}^{\mu}\right\|_{2}^{2}+\left({\sigma_{d}}^{2}+p_{d}^{\sigma}{}^{2}-2\sigma_{d}p_{d}^{\sigma}\right),$ (5)
where $\lambda_{d}$ represents the distribution distance between the current image and the stored bases for each domain $d$. Given that our training images are confined to the source and augmented domains, we define the set $\lambda$ as $\{\lambda_{1},\lambda_{2}\}$. The set $\lambda$ is processed by a softmax operation to ensure the values sum to 1, which are then used as weighting factors in the AdaIN function[19] to guide feature rectification, as detailed below:
${\mu}^{\prime}=softmax(\lambda)\cdot p^{\mu},{\sigma}^{\prime}=softmax(\lambda)\cdot p^{\sigma},$ (6)
@@
where $F$ represents the original feature map of the input image, $F_{rect}$ corresponds to its rectified feature map, $\cdot$” denotes matrix multiplication. We rectify every layer of the extracted multi-scale feature maps during both the training and inference phases to ensure a consistent workflow in the detector.
@@ III-C Object-aware Contrastive Learning Module
Object-aware gating masks. We used the bounding boxes and their corresponding classes from the annotations to generate object-aware gating masks. Specifically, let $Y=\{(b_{1},c^{1}),(b_{2},c^{2}),...,(b_{m},c^{m})\}$ represent the annotations, where $b_{m}$ and $c^{m}$ denote the bounding box and category of the $m$-th object, respectively. We first aggregate all instances of the same category to obtain the coordinate set $B^{c}=\{b^{c}_{1},b^{c}_{2},...,b^{c}_{i}\}$, where each $b$ includes two coordinates represented as $(x_{min},y_{min},x_{max},y_{max})$. Then, we generate the gating masks based on the coordinate set $B^{c}$ as follows:
@@
Instance-level feature aggregation with class queries. We insert $c$ class queries $Q=\{{q}_{1},q_{2},\ldots,q_{c}\}$ into the encoder, with each query $q_{i}\in\mathbb{R}^{d}$ being a learnable vector designed to aggregate features of a specific category using a cross-attention mechanism. The scope of aggregation is controlled by the object-aware gating masks previously obtained, enabling aggregation solely for instance-level features. Specifically, we define the feature tokens as $T$, with corresponding position encodings $P$, and object-aware gating masks as $M=\{{G}^{1},G^{2},\ldots,G^{c}\}$. In each transformer block of the encoder, all class queries $Q$ are updated in parallel through cross-attention with the feature tokens, as follows:
@@
Contrastive learning for feature alignment. Since the input images to our method include both the source and augmented domains, we can obtain class queries from two distinct domains. By attracting positive sample pairs (queries from the same category across both domains) and repelling negative samples, the detector is compelled to extract domain-invariant features from objects. Specifically, we define $Q^{S}$ and $Q^{A}\in\mathbb{R}^{C\times d}$ as class query sets from the source and augmented domains, respectively. The contrastive learning loss is constructed as follows:
@@
where $\cdot$” is used to measure the similarity between queries from different domains, and $C$ represents the total number of categories in the dataset. We calculate the loss in each block of the encoder.
@@ IV EXPERIMENTS
TABLE I: Ablation study of SA-DETR
Model Component $mAP_{50}$
Source domain Unseen target domain
Augmentation DSA OCL Daytime-Clear Daytime-Foggy Dusk-Rainy Night-Rainy Night-Clear
61.6 38.0 35.7 17.3 43.7
$\surd$ 62.6 41.1 36.0 17.6 44.8
$\surd$ 62.0 38.6 42.8 21.1 41.9
$\surd$ $\surd$ 63.8 41.7 44.2 22.2 44.7
$\surd$ $\surd$ $\surd$ 64.8 42.6 45.4 23.0 46.0
@@ IV-A Datasets
Following[4, 6], we employed a benchmark for single-source domain generalization in object detection to evaluate the SA-DETR. The dataset consists of scenes under five different weather conditions, meticulously selected from three public datasets: Berkeley Deep Drive 100K (BBD-100K)[43], Cityscapes[44], and Adverse-Weather[45]. Specifically, the experiment designated ”Daytime-Clear” as the source domain, which includes 19,395 training images and 8,313 testing images. The other four conditions, used exclusively for testing as unseen scenarios, included ”Daytime-Foggy” with 3,775 test images, ”Dusk-Rainy” with 3,501 test images, ”Night-Clear” with 26,158 test images, and ”Night-Rainy” with 2,494 test images. All datasets encompassed seven categories: person, car, bike, rider, motor, bus, and truck.
@@ IV-C Ablation Studies
In this section, we conducted ablation studies by selectively removing components to demonstrate their contributions, and the results are presented in Table I. First, we compared the performance before and after data augmentation to demonstrate that augmentation methods are suboptimal in enhancing the generalization of detection when significant scene differences are encountered. As shown in rows 1 and 2 of the table, ”Augmentation” refers to the data augmentation method[37], which demonstrates limited effectiveness in enhancing performance in scenarios such as ”Dusk-Rainy” and ”Night-Rainy”. Subsequently, the results in row 3 show that our proposed Domain Style Adapter (DSA) achieved a satisfactory performance improvement in the aforementioned challenging scenarios, effectively complementing the augmentation method. Consequently, by integrating these two methods, the generalization of the detector is substantially enhanced across all unseen scenarios, as presented in row 4. To further enhance the generalization of the detector, we propose the Object-aware Contrastive Learning (OCL) module. Unlike the proposed adapter that addresses the SDG challenge at the image level, this module improves the detector’s generalization by encouraging the extraction of domain-invariant instance-level features. As shown in row 5, the detector’s performance is further improved across various scenarios by incorporating the OCL module. Ultimately, our SA-DETR achieves significant improvements in all experimental scenarios compared to the vanilla DINO detector, demonstrating that the proposed components effectively enhance the detector’s generalization.
@@ IV-D Comparisons with the State-of-the-Art
TABLE II: Experimental results (%) on Daytime-Clear Scene.
@@
SA-DETR 64.9 54.5 85.4 54.5 69.5 58.4 66.3 64.8
TABLE III: Experimental results (%) on Dusk-Rainy Scene.
@@
SA-DETR 50.1 31.8 77.1 28.0 50.1 27.8 52.6 45.4
TABLE IV: Experimental results (%) on Night-Rainy Scene.
@@
SA-DETR 34.7 14.6 47.9 0.7 23.6 12.2 27.3 23.0
TABLE V: Experimental results (%) on Daytime-Foggy Scene.
@@
SA-DETR 39.6 33.3 67.7 34.4 48.5 47.2 27.6 42.6
TABLE VI: Experimental results (%) on Night-Clear Scene.
@@
SA-DETR 50.0 41.5 72.6 18.4 56.6 32.8 50.0 46.0
Results on Daytime-Clear Scene. We first evaluate the model’s performance on the source domain. As shown in Table II, the vanilla DINO naturally surpasses CNN-based detectors, and our method further achieves a $64.8\%$ mAP, outperforming the state-of-the-art method by $6.2\%$. This can be attributed to the strong ability of the DETR-based detector, as well as the effective data augmentation and the proposed OCL module, which further enhance the detector’s performance.
Results on Dusk-Rainy and Night-Rainy Scene. We evaluated our method’s performance in two unseen scenarios: “Dusk-Rainy” and “Night-Rainy”, with results presented in Tables III and IV. In the “Dusk-Rainy” scenario, the performance of the vanilla DINO already surpassed the best results by $2.5\%$ mAP. Our method improved upon this, achieving a $45.4\%$ mAP with an additional increase of $9.7\%$. In the Night-Rainy scenario, the vanilla DINO performs below expectations, indicating that the existing domain gap between training and testing data indeed significantly reduces the detector’s accuracy. Meanwhile, our method also demonstrated outstanding performance, achieving a $23.0\%$ mAP, which exceeds the state-of-the-art method by a margin of $3.8\%$ in mAP. This shows that the proposed DSA effectively enhances the detector’s robustness against significant distribution differences between the training and testing scenarios.
Results on Daytime-Foggy and Night-Clear Scene. Tables Vand VI show the results of methods in two other scenarios that have smaller stylistic differences compared to those discussed above. The vanilla DINO also achieved competitive results, demonstrating that DETR-based detectors have better generalization for various unseen scenarios. Our method achieves the best performance, surpassing the state-of-the-art methods by $3.0\%$ and $5.2\%$ in mAP, respectively. These improvements are primarily achieved through the combined effects of our adapter and OCL module. The experiments demonstrate SA-DETR’s powerful capability in addressing SDG object detection tasks.

### v2

@@ III-B Online Domain Style Adapter
Given the feature map of a specific layer $F\in\mathbb{R}^{(B\times C\times H\times W)}$, we first compute its statistical data, i.e., the mean $\mu$ and variance $\sigma$ along the channel dimension, as follows:
$\mu\left(F\right)=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}F,$ (1)
$\sigma\left(F\right)=\sqrt{\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}\left(F-\mu\left(F\right)\right)^{2}+\epsilon},$ (2)
where $\mu$ and $\sigma\in\mathbb{R}^{(B\times C)}$, $\epsilon$ is set to $1\times 10^{-6}$ to avoid numerical computation resulting in zero. In this manner, we obtain channel-wise statistics for the feature maps of each layer, which are subsequently leveraged to construct the style memory bank and to drive the ensuing style adaptation procedure.
Next, we design a dynamic style memory bank that stores the statistics extracted at every training iteration and supports on-the-fly style adaptation during inference. The bank is implemented as a ring buffer that maintains $K$ style prototypes and is updated after each mini batch. Concretely, the mean and deviation of every incoming feature map are compared with all stored prototypes, and their style discrepancy is measured by the Wasserstein distance, as follows:
$d_{i}\bigl{(}u,\,p_{u}^{\,i},\,\sigma,\,p_{\sigma}^{\,i}\bigr{)}=\left\lVert u-p_{u}^{\,i}\right\rVert_{2}^{2}+\bigl{(}\sigma^{2}+(p_{\sigma}^{\,i})^{2}-2\sigma\,p_{\sigma}^{\,i}\bigr{)},$ (3)
where $d_{i}$ measures the stylistic similarity between the input feature map and the $i$-th stored prototype. The terms $p_{u}^{\,i}$ and $p_{\sigma}^{\,i}$ denote the mean and variance of the $i$-th prototype in the memory bank, respectively.
To dynamically fuse similar style prototypes while retaining the most discriminative ones, thereby enabling the memory bank to self-organize a set of discriminative style bases, we introduce an adaptive threshold. $\tau$. The threshold is updated online based on the distance distribution of the current mini batch, and it is computed as follows:
$\tau\;=\;\frac{\alpha}{K}\,\sum_{i=1}^{K}d_{i}$ (4)
where $\alpha$ is a temperature coefficient that controls the sensitivity of the adaptive threshold. Let $d_{\min}=\min_{k}(d_{k})$ denote the distance between the current sample and its closest prototype. If $d_{\min}>\tau$, the sample exhibits a significant style discrepancy. In that case, we follow the queue update strategy [52] to locate the least frequently used prototype in the memory bank and replace it with the statistics of the current sample. Otherwise, we fuse the current statistics with the nearest prototype by applying an Exponential Moving Average (EMA) [53], thereby updating its mean and standard deviation to preserve the continuity and stability of the bank, as detailed below:
$p_{\mu}^{\prime}=\lambda p_{\mu}+(1-\lambda)\mu,$ (5)
$p_{\sigma}^{\prime}=\lambda p_{\sigma}+(1-\lambda)\sigma,$ (6)
where $\lambda$ denotes the momentum factor. Our memory bank follows a test-time adaptation design. During testing, we employ the same update strategy to enable the model to rapidly and continuously absorb the style characteristics of unseen scenes. It is noteworthy that only the fusion is employed to mitigate the contamination of the memory bank caused by discrete or anomalous samples during the TTA phase.
Finally, we employ the style prototypes stored in the memory bank to rectify the input image by mapping the target domain statistics to the training domain distribution. To maintain consistency with previous stages, we compute the weighting coefficients using the Wasserstein distance between the current statistics and each style prototype. Additionally, we explore a soft-KNN retrieval mechanism that selects a small set of representative bases instead of utilizing the entire bank, which is found to be suboptimal. Therefore, the resulting distances $d=\{d_{1},d_{2},...,d_{K}\}$ are normalized using softmax, so that the weights sum to 1. These normalized weights act as coefficients in AdaIN, steering the subsequent feature projection as follows:
${\mu}^{\prime}=softmax(d)\cdot p_{\mu},{\sigma}^{\prime}=softmax(d)\cdot p_{\sigma},$ (7)
@@
where $F$ represents the original feature map of the input image, $F_{rect}$ corresponds to its rectified feature map, $\cdot$” denotes matrix multiplication. We rectify every layer of the extracted multiscale feature maps during both the training and inference phases to ensure a consistent workflow in the detector.
@@ III-C Object-aware Contrastive Learning Module
III-C1 Object-aware gating masks
We used the bounding boxes and their corresponding classes from the annotations to generate object-aware gating masks. Specifically, let $Y=\{(b_{1},c^{1}),(b_{2},c^{2}),...,(b_{m},c^{m})\}$ represent the annotations, where $b_{m}$ and $c^{m}$ denote the bounding box and category of the $m$-th object, respectively. We first aggregate all instances of the same category to obtain the coordinate set $B^{c}=\{b^{c}_{1},b^{c}_{2},...,b^{c}_{i}\}$, where each $b$ includes two coordinates represented as $(x_{min},y_{min},x_{max},y_{max})$. Then, we generate the gating masks based on the coordinate set $B^{c}$ as follows:
@@ III-C1 Object-aware gating masks
III-C2 Instance-level feature aggregation with class queries
We insert $c$ class queries $Q=\{{q}_{1},q_{2},\ldots,q_{c}\}$ into the encoder, with each query $q_{i}\in\mathbb{R}^{d}$ being a learnable vector designed to aggregate features of a specific category using a cross-attention mechanism. The scope of aggregation is controlled by the object-aware gating masks previously obtained, enabling aggregation solely for instance-level features. Specifically, we define the feature tokens as $T$, with corresponding position encodings $P$, and object-aware gating masks as $M=\{{G}^{1},G^{2},\ldots,G^{c}\}$. In each transformer block of the encoder, all class queries $Q$ are updated in parallel through cross-attention with the feature tokens, as follows:
@@ III-C2 Instance-level feature aggregation with class queries
III-C3 Contrastive learning for feature alignment
Since the input images to our method include both the source and augmented domains, we can obtain class queries from two distinct domains. By attracting positive sample pairs (queries from the same category across both domains) and repelling negative samples, the detector is compelled to extract domain-invariant features from objects. Specifically, we define $Q^{S}$ and $Q^{A}\in\mathbb{R}^{C\times d}$ as class query sets from the source and augmented domains, respectively. The contrastive learning loss is constructed as follows:
@@ III-C3 Contrastive learning for feature alignment
where $\cdot$” is used to measure the similarity between queries from different domains, and $C$ represents the total number of categories in the dataset. Meanwhile, we compute the loss only for the categories that are present in the images.
@@ IV EXPERIMENTS
TABLE I: Ablation study of SA-DETR. All numbers are $mAP_{50}$ (%).
Model Component Aug. ODS‑Adapter OCL Source domain Unseen target domains
Std. TTA Day‑C Day‑F Dusk‑R Night‑R Night‑C
vanilla DINO 61.6 38.0 35.7 17.3 43.7
+Aug 63.8 41.3 37.5 18.6 44.5
+Adapter 64.2 41.7 45.1 20.7 45.0
+Adapter & TTA 64.3 42.1 45.3 22.9 45.4
SA‑DETR 64.3 41.7 46.5 24.5 45.6
@@ IV-A Datasets
Following[4, 6], we employed a benchmark for single-source domain generalization in object detection to evaluate the SA-DETR. The dataset consists of scenes under five different weather conditions, meticulously selected from three public datasets: Berkeley Deep Drive 100K (BDD-100K)[54], Cityscapes[55], and Adverse-Weather[56]. Specifically, the experiment designated ”Daytime-Clear” as the source domain, which includes 19,395 training images and 8,313 testing images. The other four conditions, used exclusively for testing as unseen scenarios, included ”Daytime-Foggy” with 3,775 test images, ”Dusk-Rainy” with 3,501 test images, ”Night-Clear” with 26,158 test images, and ”Night-Rainy” with 2,494 test images. All datasets encompassed seven categories: person, car, bike, rider, motor, bus, and truck.
@@ IV-C Ablation Studies
In this section, we first conduct ablation studies by selectively removing individual components to quantify their contributions to our framework. Next, we focus on the proposed ODS-Adapter, analyzing the effect of the number of stored style bases and the role of the constructed dynamic memory bank. Finally, we report the runtime cost of style rectification to highlight the efficiency of our approach.
IV-C1 Effectiveness of Each Component
Table I reports the ablation study, highlighting the contribution of each component. As shown in rows 1 and 2 of the table, ”Aug.” refers to the data-augmentation method [32], which yields only marginal gains in challenging conditions such as ”Dusk-Rainy” and ”Night-Rainy”. These results confirm that conventional augmentation alone is insufficient when test scenes deviate substantially from the training domain. The results in row 3 show that adding our ODS-adapter markedly improves performance on these hard domains while maintaining the accuracy of the source domain. Enabling the adapter in test-time adaptation mode further raises ”Night-Rainy” mAP to 22.9% and increases the average target domain mAP, demonstrating that online prototype updates effectively complement the offline style learned during training.
To further enhance the generalization of the detector, we propose the OCL module. This module improves the detector’s generalization by encouraging the extraction of domain-invariant instance features. As shown in row 5, the detector’s performance is further improved across various scenarios by incorporating the OCL module. However, we observe a modest performance drop in the ”Daytime-Foggy” test scenes that closely resemble the training domain, possibly because subtle domain specific cues such as slight contrast variations assist the detector in localizing objects. Ultimately, our SA-DETR achieves significant improvements in all experimental scenarios compared to the vanilla DINO detector, demonstrating that the proposed components effectively enhance the detector’s generalization.
IV-C2 Impact of the Number of Stored Style Bases
Table II investigates the effect of the number of stored style bases on detection performance. When $k=1$, all style statistics are collapsed into a single prototype, similar to existing methods [40, 9, 10], resulting in a sub-optimal average $mAP_{50}$ of 42.72%. Increasing $k$ from 1 to 4 steadily improves performance, with the largest gain appearing on the most challenging domains and a peak average of 44.52% at $k=4$. Beyond this point, adding more bases provides no additional benefit and even causes a slight decline, likely because single source training data do not offer enough stylistic diversity to populate a larger memory without redundancy or noise. These results demonstrate that the robustness of our adapter enhances the detector’s ability to adapt to unseen scenes. A modest number of bases further improves performance by striking an optimal balance between capturing meaningful intra-domain style variations and avoiding over fragmentation of the memory.
TABLE II: Impact of the number of stored style bases on $mAP_{50}$ (%).
Number Day‑C Day‑F Dusk‑R Night‑R Night‑C Avg
1 64.0 41.8 42.8 19.8 45.2 42.72
2 64.0 42.2 44.1 23.8 44.9 43.80
3 63.7 41.7 45.0 24.0 45.2 43.92
4 64.3 41.7 46.5 24.5 45.6 44.52
5 64.5 42.3 45.7 23.2 45.5 44.24
6 64.2 42.2 45.4 22.2 45.6 43.92
8 64.5 41.5 45.5 22.5 45.4 43.88
IV-C3 Hyperparameter sensitivity of the dynamic memory bank
As shown in Table III, we further validate the effectiveness of the self-organizing strategy, which discards rarely used style bases and allows the memory bank to retain a more discriminative set of style bases. “N/A” denotes using only EMA to fuse the computed statistics, which is clearly suboptimal. On the other hand, we analyze the sensitivity of the dynamic memory bank to $\alpha$ and observe only minor performance fluctuations. The best average $mAP_{50}$ is achieved when $\alpha=0.7$, largely due to more substantial gains on challenging low-light domains such as Dusk-R and Night-R.
TABLE III: Effect of the temperature coefficient $\alpha$ in the dynamic memory bank on $mAP_{50}$ (%).
$\alpha$ Day-C Day-F Dusk-R Night-R Night-C Avg.
N/A 64.4 42.0 45.2 23.4 45.3 44.06
0.6 64.7 42.3 45.1 24.3 45.9 44.46
0.7 64.3 41.7 46.5 24.5 45.6 44.52
0.8 64.4 42.5 45.4 23.9 45.3 44.30
IV-C4 Analysis of inference efficiency
Our proposed ODS-Adapter enables style rectification for unseen domains during inference and also performs dynamic updates to the style bases stored in a memory bank within a TTA framework. To evaluate the additional runtime overhead introduced during inference, we measure the average inference time over 500 runs to ensure the stability of the results, as reported in Table IV. After incorporating the Adapter, the average inference time is 87.4 ms, which introduces only a slight runtime overhead to the vanilla DINO. This modest increase in inference time can be seen as a reasonable trade-off for the performance gains brought by style rectification and dynamic style memory updates. Since we only dynamically update the style representation memory bank during testing, it not only further enhances generalization to unseen scenes but also incurs virtually no additional time overhead, demonstrating the effectiveness and efficiency of our ODS-Adapter. It is worth noting that the proposed OCL module is only used to assist the detector during training and can be removed at inference time, thus incurring no additional overhead.
TABLE IV: Average inference time per image (ms).
Method Time / ms
DINO 81.2
DINO + Adapter 87.4
DINO + Adapter & TTA 88.3
@@ IV-D Comparisons with the State-of-the-Art
TABLE V: Experimental results (%) on Daytime-Clear Scene.
@@
SA-DETR 63.8 55.8 85.1 51.9 68.9 57.2 66.4 64.3
IV-D1 Results on Daytime-Clear Scene
[…]
IV-D2 Results on Dusk-Rainy and Night-Rainy Scene
[…]
TABLE VI: Experimental results (%) on Dusk-Rainy Scene.
@@ IV-D2 Results on Dusk-Rainy and Night-Rainy Scene
SA-DETR 48.8 33.0 77.6 30.9 50.2 31.4 53.2 46.5
TABLE VII: Experimental results (%) on Night-Rainy Scene.
@@
SA-DETR 38.9 12.9 51.3 0.7 26.1 11.7 30.0 24.5
IV-D3 Results on Daytime-Foggy and Night-Clear Scene
[…]
TABLE VIII: Experimental results (%) on Daytime-Foggy Scene.
@@ IV-D3 Results on Daytime-Foggy and Night-Clear Scene
SA-DETR 38.0 31.6 67.1 31.7 48.3 46.1 28.9 41.7
TABLE IX: Experimental results (%) on Night-Clear Scene.
@@
SA-DETR 49.1 41.1 72.7 15.3 57.0 33.8 50.5 45.6